Cofab makes calculators from purchased parts. It is a manual operation consisting of three stations. Materials cost $1.25 per calculator. Several back office employees handle the orders and invoices and the company has a sales force that sells the product. The following are the cost pools and cost driver information for each activity.

Activity

Driver

Driver Frequency

Cost

Assembly

Units made

750,000

$562,500

Order processing

Orders taken

10,000

$200,000

Invoicing

Invoices mailed

25,000

$150,000

Accounting

Checks received

50,000

$250,000

Sales

Orders taken

10,000

$300,000

  1. Calculate the activity rates for each of the activity cost pools

  2. If one customer buys 5,000 units in one order, which takes 2 invoices to get payment and payment arrives in 3 checks that are processed by Accounting. What is the total cost of this customer’s order?
Answer #1

Calculation of activity rates

Assembly = $562500/750000 = 0..75

Order processing = $200000/10000 = $20

Invoicing = $150000/25000 =$6

Accounting = $250000/50000 =$5

Sales = $300000/10000 = $30

Calculation of overhead cost of order

Activity rate driver amount
Assembly $0.75 5000 $3750
Order processing $20 1 $20
Invoicing $6 2 $12
Accounting $5 3 $15
Sales $30 1 $30
Total $3827

Material cost of order = 5000 units * $1.25

= $6250

Total cost of order = $6250 + $3827 = $10077

Feel free to ask any queries..

Also plz upvote it means a lot . Thanks you